OpenStudy (anonymous):

pure algebra, find h and k |3h+4k-30|=5sqrt((h-2)^2+(k-1)^2)

OpenStudy (anonymous):

\[|3h+4k-30|=5\sqrt{(h-2)^2+(k-1)^2}\]

OpenStudy (nincompoop):

you might want to break it down first

OpenStudy (anonymous):

and?

OpenStudy (nincompoop):

and do it.

OpenStudy (anonymous):

...

OpenStudy (agent0smith):

\[\Large |3h+4k-30|=5\sqrt{(h-2)^2+(k-1)^2}\] means \[\Large 3h+4k-30=5\sqrt{(h-2)^2+(k-1)^2}\] or \[\Large 3h+4k-30= -5\sqrt{(h-2)^2+(k-1)^2}\]

OpenStudy (nincompoop):

I can give you the answer but can you figure out the solution?

OpenStudy (anonymous):

wolfram alpha gave me the answers, i want to know the steps

OpenStudy (nincompoop):

alright smith gave you the break down. start cracking

OpenStudy (anonymous):

holy fck im stupid, ok i got it wow